OpenStudy (anonymous):

At a concession stand, five hot dogs and two hamburgers cost $8.00; two hot dogs and three hamburgers cost $8.25. Find the cost of one hot dog and one hamburger.

OpenStudy (anonymous):

5d + 2h = 8 ---->(3) 2d + 3h = 8.25 -->(-2) ----------------- 15d + 6h = 24 (result of multiplying by 3) -4d - 6h = - 16.50 (result of multiplying by -2) -----------------add 11d = 7.50 d = 7.50/11 d = .68 5d + 2h = 8 5(.68) + 2h = 8 3.40 + 2h = 8 2h = 8 - 3.4 2h = 4.60 h = 4.60/2 h = 2.30 check.. 5d + 2h = 8 5(.68) + 2(2.30) = 8 3.40 + 4.60 = 8 8 = 8 (correct) hotdogs (d) cost 0.68 cents per dog hamburgers (h) cost $2.30 per burger